# DHG-Bench: A Comprehensive Benchmark on Deep Hypergraph Learning

### v1

@@ 3.3 Research Questions
We systematically design the DHG-Bench to comprehensively evaluate the existing HNN algorithms and inspire future research. In particular, we aim to investigate the following research questiones.
RQ1: How effective are the existing HNN algorithms on different benchmark tasks?
@@
Motivation and Experiment Design. Previous research on HNNs has been hindered by inconsistent data preprocessing, varying hyperparameter settings, and limited benchmark tasks, making it difficult to fairly and comprehensively evaluate the performance of different methods. Given the standardized experimental environment provided by DHG-Bench, our first research question is to revisit the effectiveness of existing HNN algorithms. By addressing this question, we aim to gain a deeper understanding of their strengths and limitations, and to identify directions for potential improvement. To this end, we evaluate the performance of HNN methods on 17 widely used hypergraph datasets across three representative tasks: node classification, hyperedge prediction, and hypergraph classification. Detailed experimental settings are provided in Appendix B, and evaluation results are presented in Section 4.1.
RQ2: How effiecient are these HNN algorithms in terms of time and space?
Motivation and Experiment Design. As existing HNN algorithms often rely on the construction of various hypergraph Laplacians or adopt a two-stage message passing scheme during training, they can naturally introduce additional computational and memory overhead compared to standard message passing GNNs. However, the efficiency and scalability of HNN methods have been largely overlooked in prior work, despite being critical for understanding the trade-off between computational cost and task performance. A thorough evaluation of these aspects is essential for assessing the practical suitability of each method for real-time and large-scale deployment. To answer this question, we evaluate the efficiency of each HNN method in terms of both time and space. Specifically, we record the total training time required to reach the best validation performance, as well as the peak GPU memory usage during training. The evaluation resutls are repo in Sec 4.2.
RQ3: How robust are the HNN algorithms under various forms of data perturbation?
Motivation and Experiment Design. Robustness evaluation is crucial for assessing the reliability of HNN algorithms under real-world conditions, especially in high-stakes domains such as healthcare and finance. Ideally, a robust HNN should maintain stable performance when exposed to various data perturbations (e.g., random noise and adversarial attacks). Such evaluation not only reveals potential vulnerabilities in existing methods but also guides the development of more resilient HNN models. To answer this research question, we conduct a systematic robustness evaluation from three perspectives: structure-level robustness, feature-level robustness, and supervision-level robustness. For structure robustness, we randomly perturb the hypergraph by adding or removing a proportion of node–hyperedge incidence connections. For feature robustness, we apply two types of random perturbations: (1) masking a certain proportion of node features, and (2) injecting Gaussian noise based on the dimensionality of node features. Furthermore, we assess the resilience of HNN algorithms to two types of supervision-level perturbations: (1) varying the sparsity of the supervision signal by randomly masking a proportion of training labels, and (2) introducing label noise by randomly assigning incorrect labels to a subset of the training samples. The evaluation results are shown in Sec 4.3.
RQ4: To what extent do existing HNN algorithms induce unfairness across demographic groups?
@@
Motivation and Experiment Design. Fairness has recently emerged as a critical concern in graph machine learning (GML) (Dong et al., 2023). Prior studies have shown that representations learned by GNNs can result in biased predictions, often favoring certain demographic groups defined by sensitive attributes (e.g., gender, race) (Ling et al., 2023; Zhu et al., 2024; Yang et al., 2024). Such bias hinders the deployment of GML models in high-stakes applications such as crime prediction (Suresh & Guttag, 2019) and credit evaluation (Yeh & Lien, 2009). Despite its importance, fairness in deep hypergraph learning has received little attention. To the best of our knowledge, this work presents the first benchmark evaluation of fairness in this context, which is crucial for developing ethically sound and reliable HNN models. To answer this question, we conduct node classification on three fairness-sensitive datasets (German, Bail, and Credit (Wei et al., 2022)), each of which contains demographic-sensitive attributes. We assess the fairness of various HNN methods using two widely adopted group fairness metrics: demographic parity () (Dwork et al., 2012), and equalized odds () (Hardt et al., 2016). Please refer to Appendix A.1 and Appendix B.2 for detailed informatoin about the datasets and metrics, respectively. The fairness evaluation results are reported in Sec 4.4.
@@ 4.1 Effectiveness Evaluation (RQ1)
| Method | Cora | Pubmed | Cora-CA | DBLP-CA | Yelp | Warmart | Trivago | Actor | Ratings | Gamers | Pokec |
| HGNN | 77.901.17 | 86.170.52 | 82.840.46 | 91.000.27 | 33.710.24 | 77.120.12 | 57.671.61 | 77.830.37 | 28.050.28 | 52.380.56 | 57.870.76 |
| HyperGCN | 78.381.63 | 87.420.42 | 81.651.58 | 89.510.18 | 29.290.55 | 68.750.56 | 42.391.25 | 81.820.39 | 27.340.72 | 51.320.72 | 57.510.54 |
| HCHA | 77.841.23 | 86.330.54 | 83.010.58 | 91.180.30 | 33.130.23 | 77.660.18 | 52.503.43 | 78.300.47 | 28.330.34 | 52.350.71 | 58.190.45 |
| LEGCN | 74.361.03 | 87.520.50 | 74.591.04 | 85.160.14 | OOM | 62.980.09 | 33.451.45 | 85.340.45 | 28.210.50 | 51.310.65 | 59.660.63 |
| HyperND | 79.230.63 | 86.730.56 | 83.190.71 | 91.340.19 | OOM | 75.100.54 | 87.191.89 | 83.190.92 | 28.320.38 | 52.390.60 | 57.651.08 |
| PhenomNN | 78.971.41 | 87.810.12 | 84.051.05 | 91.830.25 | OOM | OOM | OOM | 83.140.49 | 28.490.41 | 51.800.73 | 58.430.92 |
| SheafHyperGNN | 79.030.90 | 87.100.47 | 84.080.50 | 91.090.31 | OOM | OOM | OOM | 85.000.32 | 28.350.57 | 52.070.53 | 59.060.37 |
| HJRL | 78.671.47 | 87.980.49 | 83.720.74 | OOM | OOM | OOM | OOM | 71.540.64 | 26.900.55 | 51.620.61 | 57.570.47 |
| DPHGNN | 76.401.36 | 86.720.33 | 82.131.13 | OOM | OOM | OOM | OOM | 83.650.59 | 28.571.07 | 52.360.59 | 58.200.58 |
| TF-HNN | 79.471.31 | 87.900.37 | 84.190.89 | 91.380.24 | 35.160.54 | 77.040.12 | 90.790.79 | 85.960.41 | 28.560.68 | 52.340.53 | 59.170.52 |
| HNHN | 75.241.38 | 85.661.28 | 76.511.34 | 85.840.07 | 25.860.63 | 65.210.28 | 53.751.43 | 81.200.36 | 27.290.70 | 51.120.65 | 58.550.93 |
| UniGNN | 79.411.24 | 87.570.54 | 83.491.58 | 91.710.20 | 31.090.61 | 76.260.58 | 36.150.56 | 84.610.46 | 28.390.64 | 52.500.57 | 58.560.73 |
| AllSetTransformer | 78.021.43 | 87.790.30 | 82.950.62 | 91.510.22 | 33.180.88 | 78.610.13 | 59.924.02 | 85.660.41 | 27.321.11 | 51.740.75 | 58.550.56 |
| ED-HNN | 78.580.52 | 87.650.23 | 82.980.93 | 91.550.19 | 34.840.93 | 77.900.21 | 75.992.60 | 85.770.46 | 28.380.31 | 50.540.23 | 58.680.40 |
| EHNN | 76.511.52 | 87.120.31 | 81.680.81 | 90.470.43 | 34.093.19 | 77.950.14 | OOM | 86.210.49 | 28.950.81 | 52.140.76 | 58.231.07 |
| T-HyperGNN | 74.201.37 | 86.280.62 | 75.011.44 | 85.440.14 | OOM | 73.480.33 | OOM | 85.320.48 | 24.631.22 | 51.820.38 | 58.820.49 |
To investigate the progress and effective ranges of existing HNN methods, we compare their performance across benchmark tasks at the node, edge, and graph levels.
4.1.1 Effectiveness on node classification task. We conduct node classification experiments on 11 datasets spanning both homophilic and heterophilic hypergraphs of varying scales. As shown in Table 1, we have the following key observations: (1) TFHNN consistently ranks among the top-performing methods across diverse datasets Specifically, it achieves the best or second-best results on 8 out of 11 datasets and delivers prediction accuracy comparable to the top-performing models on the remaining three. These results demonstrate the effectiveness and robustness of its training-free message-passing module and decoupled architecture in the node classification task. (2) Existing HNN methods generally achieve lower node classification performance on heterophilic datasets than on homophilic ones. For instance, on four homophilic academic networks, all methods consistently achieve accuracies above 74%, whereas on the heterophilic Yelp and Ratings datasets, the highest accuracy is below 36%. The results indicate that heterophilic patterns can impair the quality of hypergraph representation learning. (3) Most existing HNN algorithms either face scalability issues or experience significant accuracy degradation on large-scale datasets. On the Trivago dataset, for example, six methods encounter OOM errors, and eight of the remaining ten achieve accuracies below 80%, with at least a 14.8% gap from TFHNN, the best-performing model on this dataset.
4.1.2 Effectiveness on hyperedge prediction task. For the edge-level task, we conduct experiments on six widely used real-world datasets, reporting AUROC and AP (Average Precision). As shown in Table 2, we find that: (1) Recent proposed HNN methods that achieve superior node classification performance fail to remain equally competitive in hyperedge prediction. In particular, we find that HGNN, HyperGCN, and EHNN, all proposed before 2022, together account for the best and second-best performance on the six datasets for the edge-level task. In contrast, more recent HNNs (e.g., HJRL, TFHNN, and DPHGNN) often show a notable performance gap compared to these three. For example, on Cora-CA, DPHGNN achieves an AUROC of 82.89% and an AP of 83.78%, representing decreases of 9.84% and 9.64%, respectively, relative to HyperGCN. (2) The performance gaps between most HNN algorithms and the corresponding state-of-the-art models vary greatly across datasets, with the magnitude of the gap heavily dependent on the dataset. For instance, on the Pubmed dataset, PhenomNN, UniGNN, and AllSetTransformer each achieve an AUROC within 5.5% of the state-of-the-art EHNN (79.54%). In contrast, on the Actor dataset, their AUROC scores fall at least 15% below that of the best-performing model, HGNN. This suggests that current research lacks HNN methods with strong generalization ability for hyperedge prediction across diverse datasets.
4.1.3 Effectiveness on hypergraph classification task. We conduct hypergraph classification task on two synthetic datasets and four real-world datasets. As current tensor-based methods cannot be extend to graph-level tasks, we do not report their results. Based on the results presented in Table 3, we summarize the following key findings: (1) All HNN methods achieve a substantial performance advantage on synthetic hypergraph classification datasets over real-world ones. The inferior performance on real-world datasets reflects the higher structural complexity of real-world hypergraphs, underscoring the importance of designing and evaluating algorithms on real-world data. (2) The classification performance of HNN algorithms varies greatly across datasets, with no method consistently achieving state-of-the-art results. For example, while DPHGNN achieves the highest accuracy on IMDB-Dir-Form, it ranks only 11th among 14 methods on IMDB-Dir-Genre, suggesting that dataset characteristics can significantly affect model performance. (3) Higher accuracy does not necessarily guarantee improved macro-F1 performance, particularly in class-imbalanced settings. For example, on the imbalanced Twitter-Friend dataset, HNHN achieves the third-highest accuracy among all HNN methods, yet obtains a macro-F1 of only 39.40%, which is the lowest among all methods. Conversely, HyperND attains the best macro-F1 on the same dataset but ranks only 10th in accuracy out of the 14 evaluated models.
| Method | Cora | PubMed | Cora-CA | DBLP-CA | Actor | Pokec | ||||||
| AUROC | AP | AUROC | AP | AUROC | AP | AUROC | AP | AUROC | AP | AUROC | AP | |
| SheafHyperGNN | 96.001.38 | 95.961.32 | 99.730.33 | 99.740.32 | 62.342.06 | 56.473.49 | 77.001.14 | 72.781.17 | 53.112.39 | 52.562.74 | 56.492.51 | 51.434.42 |
| HJRL | 96.100.80 | 95.980.85 | 99.600.53 | 99.600.53 | 63.092.83 | 56.543.62 | 77.821.47 | 73.731.92 | 51.843.52 | 51.133.29 | 57.102.79 | 44.197.19 |
@@ 4.2 Efficiency and Scalability Evaluation (RQ2)
To answer RQ2, we analyze the training time and memory consumption of all methods on datasets of varying scales, including Cora, DBLP-CA, Yelp, and Trivago. From Figure 2, we make the following observations: (1) Most HNN methods struggle to achieve a satisfactory balance between predictive performance and efficiency. For example, on the Yelp dataset, although EDHNN and EHNN achieve higher accuracy than the simple HGNN method, their training times are more than 9x and 23x longer, respectively. Similarly, many HNNs consume excessive memory, particularly on large-scale datasets. On Yelp, 7 out of 16 methods encounter out-of-memory (OOM) issues. On the Trivago dataset, although several HNNs remain computationally scalable, most methods fail to deliver satisfactory predictive performance. Only 2 out of 10 methods, TFHNN (90.79%) and HyperND (87.19%), achieve accuracy above 60%. (2) Compared to spectral-based and spatial-based methods, tensor-based approaches generally suffer from more severe efficiency and scalability challenges. T-HyperGNN can only scale to the medium-sized DBLP-CA dataset, where it runs approximately 406 times slower than the fastest HGNN method. Likewise, EHNN is the slowest method on Yelp and fails to scale to the large-scale Trivago dataset. (3) Among all evaluated methods, TFHNN demonstrates the most favorable trade-off between effectiveness, efficiency, and scalability. In particular, it achieves the highest accuracy on three out of the four datasets. On the large-scale Trivago dataset, it delivers the best predictive performance while consuming no more than 1.6 GB of memory and requiring less than 30 seconds of runtime, ranking first in memory usage and second in training time among all methods. These results highlight the strong potential of the decoupled architecture for designing efficient and scalable HNN algorithms. Building on these findings, future research should place greater emphasis on innovative architectural design to achieve a better balance among effectiveness, efficiency, and scalability, thereby facilitating the development of more practical HNN solutions.

### v2

@@ 3.3 Research Questions
We systematically design the DHG-Bench to comprehensively evaluate the existing HNN algorithms and inspire future research. In particular, we aim to investigate the following research questions.
RQ1: How much progress has been made by existing HNN methods?
@@
Motivation and Experiment Design. Previous research on HNNs has been limited by inconsistent experimental settings and insufficient coverage of datasets, algorithms, and tasks, thereby hindering fair and comprehensive evaluation of different methods. Given the standardized experimental environment provided by DHG-Bench, the first question is to revisit the progress of existing HNN methods and identify potential directions for enhancement. A high-quality HNN method is expected to perform consistently well across different datasets and application scenarios. To answer this question, we evaluate the performance of HNN methods on diverse, widely used hypergraph datasets across three benchmark tasks: node classification, hyperedge prediction, and hypergraph classification. Detailed experimental settings can be found in Appendix B.1.
RQ2: How efficient are these HNN methods in terms of time and space?
Motivation and Experiment Design. Training the message-passing module of HNNs makes loss computation interdependent for connected nodes, resulting in intensive computational demands and substantial memory constraints. However, the efficiency and scalability of HNN algorithms have been largely overlooked. A thorough understanding of the trade-off between computational cost and predictive performance is essential for assessing their suitability for real-time and large-scale applications. To answer this question, we perform node classification, the most widely used benchmark task, on datasets of varying scales (Cora, DBLP-CA, Yelp, and Trivago), reporting the training time to reach the best validation performance and the peak GPU memory consumption.
RQ3: Are existing HNN methods robust to different types of data perturbations?
Motivation and Experiment Design. Real-world hypergraph data inevitably contains noise, task-irrelevant information, or even mistakes (Cai et al., 2022). A reliable HNN should maintain stable performance when exposed to such noisy data, particularly in high-stakes domains such as healthcare and finance (Cai et al., 2025), where inaccurate decisions can adversely affect individual lives or broader societal systems. Evaluating the robustness of HNNs not only reveals potential vulnerabilities in existing methods but also guides the development of more resilient models. To answer this question, we simulate realistic data perturbations from three perspectives: structure, feature, and supervision signals. For each perturbation type, we vary the noise intensity and subsequently train and test HNNs on the corresponding modified hypergraph. Detailed experimental settings are in Appendix B.4.
RQ4: Do existing HNN methods yield unbiased predictions across demographic groups?
@@
Motivation and Experiment Design. Fairness has recently emerged as a critical concern in graph machine learning (GML) (Dong et al., 2023). Prior studies have shown that representations learned by GNNs can result in biased predictions, often favoring certain demographic groups defined by sensitive attributes (e.g., gender and race) (Ling et al., 2023; Zhu et al., 2024; Yang et al., 2024). Such bias hinders the deployment of GML models in high-stakes applications such as crime prediction (Suresh & Guttag, 2019) and credit evaluation (Yeh & Lien, 2009). Despite its importance, fairness in deep hypergraph learning has received little attention. To the best of our knowledge, this work presents the first benchmark evaluation of fairness in this context, which is crucial for developing ethically sound and trustworthy HNN models. To answer this question, we conduct node classification on three fairness-sensitive datasets (German, Bail, and Credit (Wei et al., 2022)), each of which contains demographic-sensitive attributes. We assess algorithmic fairness using two widely adopted group fairness metrics: demographic parity () (Dwork et al., 2012), and equalized odds () (Hardt et al., 2016). The detailed descriptions of the two metrics can be found in Appendix B.5.
@@ 4.1 Effectiveness Evaluation (RQ1)
| Method | Cora | Pubmed | Cora-CA | DBLP-CA | Walmart | Trivago | Actor | Gamers | Pokec | Yelp |
| MLP | 75.330.88 | 86.620.26 | 75.571.08 | 85.540.15 | 63.210.12 | 36.760.66 | 86.060.36 | 52.570.49 | 59.640.48 | 31.840.45 |
| CEGCN | 76.900.75 | 86.030.39 | 78.401.25 | 89.750.33 | 70.400.18 | 47.241.09 | 67.410.29 | 51.020.53 | 57.370.38 | OOM |
| CEGAT | 77.221.03 | 86.090.51 | 78.021.24 | 89.610.22 | 65.830.92 | OOM | 73.870.83 | 51.050.61 | 57.340.52 | OOM |
| HGNN | 77.901.17 | 86.170.52 | 82.840.46 | 91.000.27 | 77.120.12 | 57.671.61 | 77.830.37 | 52.380.56 | 57.870.76 | 33.710.24 |
| HyperGCN | 78.381.63 | 87.420.42 | 81.651.58 | 89.510.18 | 68.750.56 | 42.391.25 | 81.820.39 | 51.320.72 | 57.510.54 | 29.290.55 |
| HCHA | 77.841.23 | 86.330.54 | 83.010.58 | 91.180.30 | 77.660.18 | 52.503.43 | 78.300.47 | 52.350.71 | 58.190.45 | 33.130.23 |
| LEGCN | 74.361.03 | 87.520.50 | 74.591.04 | 85.160.14 | 62.980.09 | 33.451.45 | 85.340.45 | 51.310.65 | 59.660.63 | OOM |
| HyperND | 79.230.63 | 86.730.56 | 83.190.71 | 91.340.19 | 75.100.54 | 87.191.89 | 83.190.92 | 52.390.60 | 57.651.08 | OOM |
| PhenomNN | 78.971.41 | 87.810.12 | 84.051.05 | 91.830.25 | OOM | OOM | 83.140.49 | 51.800.73 | 58.430.92 | OOM |
| SheafHyperGNN | 79.030.90 | 87.100.47 | 84.080.50 | 91.090.31 | OOM | OOM | 85.000.32 | 52.070.53 | 59.060.37 | OOM |
| HJRL | 78.671.47 | 87.980.49 | 83.720.74 | OOM | OOM | OOM | 71.540.64 | 51.620.61 | 57.570.47 | OOM |
| DPHGNN | 76.401.36 | 86.720.33 | 82.131.13 | OOM | OOM | OOM | 83.650.59 | 52.360.59 | 58.200.58 | OOM |
| TF-HNN | 79.471.31 | 87.900.37 | 84.190.89 | 91.380.24 | 77.040.12 | 90.790.79 | 85.960.41 | 52.340.53 | 59.170.52 | 35.160.54 |
| HNHN | 75.241.38 | 85.661.28 | 76.511.34 | 85.840.07 | 65.210.28 | 53.751.43 | 81.200.36 | 51.120.65 | 58.550.93 | 25.860.63 |
| UniGNN | 79.411.24 | 87.570.54 | 83.491.58 | 91.710.20 | 76.260.58 | 36.150.56 | 84.610.46 | 52.500.57 | 58.560.73 | 31.090.61 |
| AllSetTransformer | 78.021.43 | 87.790.30 | 82.950.62 | 91.510.22 | 78.610.13 | 59.924.02 | 85.660.41 | 51.740.75 | 58.550.56 | 33.180.88 |
| ED-HNN | 78.580.52 | 87.650.23 | 82.980.93 | 91.550.19 | 77.900.21 | 75.992.60 | 85.770.46 | 50.540.23 | 58.680.40 | 34.840.93 |
| HyperGT | 75.571.11 | 86.060.54 | 75.420.62 | 84.530.30 | OOM | OOM | 84.430.47 | 51.190.57 | 57.730.76 | OOM |
| EHNN | 76.511.52 | 87.120.31 | 81.680.81 | 90.470.43 | 77.950.14 | OOM | 86.210.49 | 52.140.76 | 58.231.07 | 34.093.19 |
| T-HyperGNN | 74.201.37 | 86.280.62 | 75.011.44 | 85.440.14 | 73.480.33 | OOM | 85.320.48 | 51.820.38 | 58.820.49 | OOM |
To investigate the effectiveness of existing HNNs, we compare their performance across benchmark tasks at the node, edge, and graph levels. Due to space constraints, additional node classification results on NTU2012, ModelNet40, and Ratings (Table A5), as well as the complete results of hyperedge prediction (Table A6) and hypergraph classification (Table A7), are available in Appendix C.1.
4.1.1 Effectiveness on Node Classification Task
Results (Table 1 and Table A5). Across diverse datasets, HNNs generally outperform both CEGCN and CEGAT, suggesting that naively extending GNNs to hypergraphs via clique expansion disrupts high-order structures and degrades predictive performance. This highlights the necessity of designing neural architectures with dedicated high-order message passing. HNNs achieve notable improvements over MLP on homophilic datasets, but on heterophilic datasets, most HNNs even underperform MLP, which only leverages node features. This reveals the adverse impact of heterophilic connections on hypergraph representation learning and underscores the need to rethink HNN design in such settings. TF-HNN consistently ranks among the top-performing methods across diverse datasets, achieving optimal or near-optimal results. Moreover, unlike other advanced HNNs (e.g., PheomNN, DPHGNN, and HyperGT) that fail on large-scale datasets due to out-of-memory issues, TF-HNN remains scalable. These findings underscore the promise of its decoupled architecture for enhanced generalization and scalability.
4.1.2 Effectiveness on Hyperedge Prediction Task
Results (Table A6). Advanced HNN methods that generally achieve superior performance on node classification fail to maintain the same level of competitiveness in hyperedge prediction. Specifically, the two earliest methods, HGNN and HyperGCN, along with the tensor-based EHNN introduced in 2022, collectively achieve all the best results and the majority of second-best results across the six hyperedge prediction datasets. In contrast, recent HNNs (e.g., ED-HNN, HJRL, DPHGNN, TF-HNN) often show a notable performance gap compared to the above three. For example, on DBLP-CA, TF-HNN achieves an AUROC of 75.70% and an AP of 74.97%, which are 13.76% and 16.70% lower than those of the best-performing model, HyperGCN. Across hyperedge prediction benchmarks, HNN algorithms display considerable performance divergence depending on the dataset, and none consistently deliver the best results. For instance, while EHNN achieves state-of-the-art performance on Cora and Pubmed, it obtains only 77.83% AUROC on Cora-CA, ranking 11th among 17 HNNs and 14.90% lower than the top-performing HyperGCN.
4.1.3 Effectiveness on Hypergraph Classification Task
Results (Table A7). HNN algorithms perform markedly better on synthetic datasets than on real-world ones. On RHG-10, most models achieve over 90% accuracy and Macro-F1, and on RHG-3, many even exceed 98%. In contrast, on real-world datasets, accuracies rarely surpass 70%, reflecting the structural complexity of real hypergraphs. This gap underscores the need for more realistic and challenging benchmarks to rigorously evaluate hypergraph classification. HNN methods generally outperform GNN-based approaches built on clique expansion, as the latter often distorts global hypergraph structures, whereas higher-order message passing in HNNs preserves these dependencies and enhances discriminative power. HNNs’ performance varies considerably across datasets, with no method demonstrating consistent superiority. For instance, while DPHGNN achieves the best accuracy on IMDB-Dir-Form, it falls to 11th on IMDB-Dir-Genre and 14th on Steam-Player across all evaluated HNNs, underscoring the substantial impact of dataset characteristics on model performance. Many HNN methods fail to achieve a desirable trade-off between accuracy and Macro-F1. For example, on the Twitter dataset, HNHN achieves 58.47% accuracy (third highest among all HNN models) but only 39.40% Macro-F1, the lowest overall.
@@ 4.2 Efficiency and Scalability Evaluation (RQ2)
Results (Figure 1). CEGCN and CEGAT face scalability challenges on large datasets (e.g., Yelp and Trivago), where clique expansion produces dense edges and leads to significant training memory overhead. Most advanced HNN methods struggle to achieve a satisfactory balance between model utility and efficiency. For example, on the Yelp dataset, ED-HNN and EHNN provide only marginal accuracy gains over the simple HGNN, yet their training times are over 9× and 23× longer, respectively, reflecting a substantial rise in computational cost. In addition, many HNNs suffer from memory bottlenecks on large-scale datasets. Specifically, on Yelp, 8 out of 17 methods encounter out-of-memory (OOM) issues. On Trivago, although 10 HNNs remain computationally scalable, most fail to deliver satisfactory predictive performance. Only TF-HNN (90.79%) and HyperND (87.19%) achieve accuracy above 60%. This may result from the intricate patterns of large-scale graphs. Tensor-based approaches exhibit more pronounced efficiency and scalability limitations than the other two kinds of methods. T-HyperGNN can only scale to the medium-sized DBLP-CA dataset, where it runs approximately 406 times slower than the fastest method, HGNN. Moreover, on Yelp, EHNN incurs the longest training time and fails to scale to the large-scale Trivago dataset. Among all evaluated methods, TF-HNN generally achieves a superior trade-off between utility and both time and space efficiency. For example, on the large-scale Trivago dataset, it achieves the best predictive performance with no more than 1.6 GB of memory and under 30 seconds of runtime, ranking first in memory efficiency and second in training time among all HNN methods.